OpenStudy (anonymous):

Jessica has a compost pit in her garden. The pit is 8 feet long, 4 feet wide, and 2 feet deep. She plans to make it 1.5 feet deeper. - What will be the increase in the volume of the pit?

OpenStudy (ghazi):

initial volume = 8*4*2 and final vol = 8*4*(2+1.5) now calculate and check out the difference

OpenStudy (anonymous):

so 48

OpenStudy (anonymous):

inicial volume = 8*4*2 new volume = 8*4*(2+1.5) difference = 8*4*(2+1.5) - 8*4*2 = 8*4 ( 2+1.5 -2) = 8*4*1.5

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

Thanks guys :)

OpenStudy (ghazi):

anytime